HOW TO FIND COORDINATES OF A HELIX FORMED BY THE DEVIATION OF A PARTICLE IN A MAGNETIC FIELD AT A CERTAIN TIME?

Grade:

1 Answers

Aman Bansal
12 years ago

Dear Dhruba,

The force on a current carrying wire is similar to that of a moving charge as expected since a charge carrying wire is a collection of moving charges. A current carrying wire feels a sideways force in the presence of a magnetic field. The Lorentz force on a macroscopic current is often referred to as the Laplace force. Consider a conductor of length l and area of cross section A and has charge q which is due to electric current i .If a conductor is placed in a magnetic field of induction B which makes an angle θ (theta) with the velocity of charges in the conductor which has i current flowing in it. then force exerted due to small particle q is

\mathbf{F} = q \mathbf{v}\sin(\theta),

then for n number of charges it has

N = nlA

then force exerted on the body is

f=\mathbf{F}N=q \mathbf{v}\sin(\theta)(nlA)

but

nqvA = i

that is

f=\mathbf{B}il\sin(\theta)


Best Of luck

Cracking IIT just got more exciting,It s not just all about getting assistance from IITians, alongside Target Achievement and Rewards play an important role. ASKIITIANS has it all for you, wherein you get assistance only from IITians for your preparation and win by answering queries in the discussion forums. Reward points 5 + 15 for all those who upload their pic and download the ASKIITIANS Toolbar, just a simple  to download the toolbar….

So start the brain storming…. become a leader with Elite Expert League ASKIITIANS

Thanks

Aman Bansal

Askiitian Expert